OpenStudy (anonymous):

the annual population of a country is increasing at a rate of 0.69% that is dp/dt =0.069. if the population is 50000 in 1995 find P=50000 e^(0.069*t) a) in 2000 population is when t=5 Then P= 50000e^(0.069*5) which = 70599 Part b asks "the rate at which the population will be growing in the year 2000" ? Hw am i suppose to find rate ? I try finding k but it came o.o69

OpenStudy (dan815):

that is not the real differntial equation

OpenStudy (dan815):

it should be dp/dt=0.069*p

OpenStudy (anonymous):

No its application of calculus to the physical world where k constant represents as dp/dt

OpenStudy (dan815):

you can do dp/dt at t=5

OpenStudy (anonymous):

Oh wait yes i forgot to write P srry

OpenStudy (dan815):

yeah i know :P

OpenStudy (dan815):

I model equations for a living

OpenStudy (anonymous):

So 5=o0.069 P

OpenStudy (dan815):

no

OpenStudy (dan815):

look at the p(t) equation

OpenStudy (anonymous):

0.069*

OpenStudy (anonymous):

Yea so dP/ 5 = 0.069P

OpenStudy (dan815):

P(t)=50000 e^(0.069*t) P'(t)=0.069*50000*e^0.069t P'(5)=0.069*50000*e^(0.069*5)=?

OpenStudy (anonymous):

Oh i get it. Lol thnx